### v1

# A Study on Zero-shot Non-intrusive Speech Assessment using Large Language Models

Abstract
This work investigates two strategies for zero-shot non-intrusive speech assessment leveraging large language models. First, we explore the audio analysis capabilities of GPT-4o. Second, we propose GPT-Whisper, which uses Whisper as an audio-to-text module and evaluates the text’s naturalness via targeted prompt engineering. We evaluate assessment metrics predicted by GPT-4o and GPT-Whisper examining their correlations with human-based quality and intelligibility assessments, and character error rate (CER) of automatic speech recognition. Experimental results show that GPT-4o alone is not effective for audio analysis; whereas, GPT-Whisper demonstrates higher prediction, showing moderate correlation with speech quality and intelligibility, and high correlation with CER. Compared to supervised non-intrusive neural speech assessment models, namely MOS-SSL and MTI-Net, GPT-Whisper yields a notably higher Spearman’s rank correlation with Whisper’s CER. These findings validate GPT-Whisper as a reliable method for accurate zero-shot speech assessment without requiring additional training data (speech data and corresponding assessment scores).
Index Terms:
speech assessment, zero-shot, non-intrusive, whisper, ChatGPT, large language model
I Introduction
Speech assessment metrics play a critical role in evaluating a variety of speech-related applications, including speech enhancement [1, 2], hearing aid (HA) devices [3, 4, 5], and telecommunications [6]. With the advancement of deep learning and the need for accurate non-intrusive speech assessment metrics, researchers have increasingly adopted deep learning models for speech assessment [7, 8, 9, 10, 11, 12, 13]. To achieve accurate automatic assessments, various strategies have been explored, such as reducing listener bias [14], integrating large pre-trained models (e.g., self-supervised learning (SSL) models [15, 16], Whisper [11], and speech language models[17]), utilizing ensemble learning [7, 8, 9], and incorporating pseudo labels [18]. Despite significant performance improvements, achieving satisfactory generalization with limited training samples remains challenging.
Recently, with the development of large-scale conversational agents like ChatGPT, there has been an increasing interest in evaluating their ability to perform more extensive reasoning and understanding [19]. The capabilities of GPT-4, especially its latest extension GPT-4o  [^1] https://platform.openai.com/docs/models/gpt-4o, have been significantly expanded to enable not only advanced text understanding but also multimodal integration, such as merging visual and audio understanding. In the case of image understanding, ChatGPT has been successfully integrated with visual language models to perform tasks such as deep fake detection [20]. For audio understanding, a noteworthy integration is AudioGPT [21]. This approach integrates ChatGPT with multiple pre-trained audio models. Based on specific prompt input, the most appropriate audio model is selected to generate the response. While AudioGPT excels at integrating audio understanding with prompt engineering, it does not cover speech assessment. Given the interest in reliable non-intrusive speech assessment with minimal training samples, we intend to investigate whether ChatGPT can effectively perform speech assessment in a zero-shot setting and identify optimal strategies to ensure accurate and unbiased results.
In this paper, we explore two strategies for zero-shot non-intrusive speech assessment by leveraging large language models (LLM). First, we directly leverage the audio analysis capabilities of GPT-4o for speech assessment. Second, we propose a more advanced approach, namely GPT-Whisper. Specifically, we use Whisper [22] as an audio-to-text module to create text representations, which are then assessed using targeted prompt engineering focused on evaluating the naturalness of the predicted text. To the best of our knowledge, this is the first attempt to leverage ChatGPT for speech assessment. Furthermore, to gain a deeper insight into our new assessment metric computed by GPT-Whisper, we also evaluate its correlation with several metrics, including human-based quality assessment, human-based intelligibility assessment, and the character error rate (CER) of automatic speech recognition (ASR) models. We use two open-source ASR models: Whisper [22], and Google ASR [23] to generate CER scores. Finally, we compare our approach with two supervised neural speech assessment models: MOS-SSL [15] and MTI-Net [24].
Our experimental results first confirm that using ChatGPT alone for audio analysis is not optimal because it mainly relies on factors such as amplitude range, standard deviation, and signal-to-noise ratio (SNR) to assess the quality or intelligibility of audio data. Next, in terms of Spearman’s rank correlation coefficient (SRCC) [25], our GPT-Whisper metric yields moderate correlation with human-based quality assessment (0.4360), human-based intelligibility assessment (0.5485), and Google ASR’s CER (0.5415) and high correlation with Whisper’s CER (0.7784). This further validates our new metric as a viable assessment for evaluating speech in a zero-shot manner, especially given its higher correlation with CER. Finally, when comparing GPT-Whisper with MOS-SSL and all variants of MTI-Net (wav2vec and Whisper) in predicting Whisper’s CER, experimental results show that GPT-Whisper outperforms them in terms of SRCC (0.7784 vs. 0.7482, 0.7418, and 0.7655).
The remainder of this paper is organized as follows. Section II presents the proposed methodology. Section III describes the experimental setup and results. Finally, Section IV presents the conclusions and future work.
Figure 1: Zero-shot speech assessment with GPT-4o.
II Methodology
II-A GPT-4o for speech assessment
The main intention behind using GPT-4o for speech assessment is to validate ChatGPT’s reasoning capabilities in understanding speech characteristics. The overall framework for zero-shot speech assessment using GPT-4o is shown in Fig. 1. Specifically, given an audio input A and a prompt $P$, ChatGPT utilizes this information to estimate the assessment metric $S_{GPT-4o}$, as follows:
$\small\begin{array}[]{c}S_{GPT-4o}=ChatGPT(A,P).\end{array}$ (1)
Figure 2: Zero-shot speech assessment with GPT-Whisper.
II-B GPT-Whisper
The overall framework of GPT-Whisper is shown in Fig. 2. The main idea behind GPT-Whisper is to leverage the reasoning capabilities of GPT-4o to assess the predicted word sequence T from the audio input A. We specifically chose Whisper [22] as the audio-to-text module because of its outstanding capability to convert audio into text across a variety of acoustic environments. We assume that this capability enables Whisper to capture subtle nuances of spoken language. Specifically, given an input audio A, the audio-to-text conversion using Whisper ASR is defined as follows:
$\small\begin{array}[]{c}T=Whisper(A).\end{array}$ (2)
Based on the predicted word sequence T, we use ChatGPT to estimate the GPT-Whisper score $S_{GPT-Whisper}$. Our strategy involves using naturalness in prompt engineering, which refers to how similar the predicted text is to human-generated text in terms of fluency, coherence, and context. This approach measures the degree to which text reflects the natural flow and nuances of human speech. The process of estimating the GPT-Whisper score is defined as follows:
$\small\begin{array}[]{c}S_{GPT-Whisper}=ChatGPT(T,P).\end{array}$ (3)
III Experiments
III-A Experimental setup
The two proposed approaches are evaluated on the TMHINT-QI(S) dataset [11], which is an extension of the TMHINT-QI dataset [26], by including additional unseen noises, speakers, and enhancement systems. The TMHINT-QI(S) dataset is also a benchmark track in the VoiceMOS Challenge 2023 [27]. The evaluation set consists of noisy, clean, and enhanced utterances, covering three seen noise conditions (babble, white, and pink) and one unseen noise condition (street). It also includes three seen speech enhancement systems (Minimum-Mean Square Error (MMSE) [28], Fully Convolutional Network (FCN) [29], and Transformer [30]), and introduces two new, unseen speech enhancement systems: Conformer-based Metric Generative Adversarial Network (CMGAN) [31] and DEMUCS [32]. In total, there are 1,960 utterances with corresponding quality scores (0-5) and intelligibility scores (0-1). An additional metric, CER (0-1), is produced by two ASR models (Google ASR [23] and Whisper [22]). Note that we reverse the CER score so that higher values indicate better performance, ensuring consistency with other metrics and simplifying interpretation.
Two evaluation metrics, linear correlation coefficient (LCC) and Spearman’s rank correlation coefficient (SRCC) [25], are used to measure the performance of the proposed methods. Higher LCC and SRCC values indicate a stronger correlation between the predicted score and the ground-truth score, indicating better assessment prediction performance.
Figure 3: Example of GPT-4o used for speech assessment.
III-B Experimental results of GPT-4o
In the first experiment, we intend to leverage prompt engineering on GPT-4o for speech assessment. As shown in Fig. 3, we specifically use keywords to evaluate the quality and intelligibility of a given audio input. The audio input in Fig. 3 is an MMSE-enhanced speech of noisy speech containing babble noise at -2dB SNR, which is used as a representative scenario. The results show that GPT-4o determines the quality or intelligibility based only on metrics such as SNR, clipping, amplitude, or loudness. However, when we compare these results to the corresponding ground-truth information (quality score: 1.4/5, intelligibility score: 0.6/1, SNR: -2dB), GPT-4o makes inaccurate predictions. For example, as shown in Fig. 3, GPT-4o believes that the audio input is unclipped and has a high SNR, whereas in fact the audio has a low SNR (MMSE’s noise reduction effect is noticeably poor in this example) and moderate distortion. Furthermore, while the intelligibility estimate is fairly close to the true value (7/10 vs 0.6/1), the quality estimate is quite off (7/10 vs 1.4/5). Experimental results on the complete test set covering different audio characteristics show that it is difficult for the current GPT-4o to precisely predict the quality and intelligibility of the audio input. Therefore, the current GPT-4o may not serve as a reliable zero-shot non-intrusive speech assessment method.
III-C Experimental results of GPT-Whisper
In this section, we aim to evaluate the capabilities of our proposed GPT-Whisper for performing zero-shot non-intrusive speech assessment.
III-C1 Correlation analysis
In initial experiments, we estimate GPT-Whisper scores using the approach outlined in Section II. Experimental results show that GPT-4o exhibits notable reasoning capabilities in text comprehension and demonstrates solid contextual understanding, as shown in Fig. 4. GPT-4o effectively evaluates the coherence of text inputs and determines whether connections are natural or unnatural. Based on these results, we conclude that the scores produced by GPT-Whisper are more reliable than those produced previously by GPT-4o for speech assessment.
Based on the promising performance of GPT-Whisper, we intend to further investigate which assessment metrics are more correlated with the GPT-Whisper score. We assume that GPT-Whisper can be used as an alternative in scenarios where there is no corresponding label, and the training data is insufficient to establish a deep learning-based speech assessment model. In our evaluation, we compare GPT-Whisper scores with human-based quality, human-based intelligibility scores, and the CER of two ASR models (Whisper and Google ASR). The results in Fig. 5 show that the GPT-Whisper metric has a higher correlation with intelligibility-based metrics than the quality one and has the highest correlation with the CER of Whisper ASR. These results further confirm GPT-Whisper’s ability to perform zero-shot speech assessment and serve as a surrogate metric for evaluating speech recognition performance, given its high correlation score with the CER of Whisper (SRCC of 0.7784).
Figure 4: Example of GPT-Whisper used for speech assessment.
Figure 5: Correlation of GPT-Whisper scores with other metrics.
III-C2 Comparison with supervised methods
We now compare GPT-Whisper with two supervised speech assessment models, namely MOS-SSL [15], and MTI-Net [24], on the task of predicting the CER of Whisper. We use Whisper’s CER as the ground truth because the GPT-Whisper score has a higher correlation with this CER score than with human-based quality and intelligibility. Therefore, it is interesting to compare GPT-Whisper with these supervised models on this metric. Specifically, MOS-SSL incorporates wav2vec 2.0 to predict MOS scores, while MTI-Net employs cross-domain features (raw waveform, spectral features, and SSL features) to predict intelligibility metrics (intelligibility, CER, and STOI) using multi-task learning [24]. We trained MOS-SSL and MTI-Net using the TMHINT-QI(s) training set. MOS-SSL was trained to predict Whisper’s CER, while MTI-Net was trained to predict intelligibility, Google ASR’s CER, Whisper’s CER, and STOI. Given the significant efficacy of integrating Whisper representations into robust speech assessment models, [11, 8, 9], we also integrated Whisper into the MTI-Net model. Therefore, we prepared two versions: MTI-Net (wav2vec) and MTI-Net (Whisper).
The results in Table I confirm that GPT-Whisper performs comparably to supervised non-intrusive speech assessment models. Interestingly, GPT-Whisper outperforms MOS-SSL and MTI-Net (wav2vec) in LCC and SRCC. Additionally, GPT-Whisper has higher SRCC than MTI-Net (Whisper), although its LCC is slightly lower. These results demonstrate the effectiveness of GPT-Whisper as a zero-shot speech assessment model, showing comparable or even superior performance to supervised methods.
TABLE I: Comparison of GPT-Whisper with supervised methods.
System Supervised LCC SRCC
GPT-Whisper No 0.7541 0.7784
MOS-SSL [15] Yes 0.7328 0.7482
MTI-Net (wav2vec) [24] Yes 0.7344 0.7418
MTI-Net (Whisper) [24] Yes 0.8194 0.7655
Figure 6: Distribution of predicted CER scores of GPT-Whisper and MTI-Net (Whisper) and Whisper’s ground-truth CER scores.
III-C3 Score distribution analysis
In this section, we further analyze the distribution of scores evaluated by GPT-Whisper and MTI-Net (Whisper) and the corresponding CER labels of Whisper. This analysis aims to evaluate whether GPT-Whisper provides accurate score estimations across the entire score range. As shown in Fig. 6, GPT-Whisper successfully predicts the entire range of scores and shows a more similar distribution to that of CER generated by the Whisper ASR, especially for CER less than 0.25. The closeness of the distribution of predicted CER scores to the distribution of true CER scores validates GPT-Whisper as a new assessment metric and the potential of using ChatGPT for speech assessment by implementing the GPT-Whisper method.
IV Conclusions
In this paper, we have proposed two strategies for zero-shot non-intrusive speech assessment. The first strategy leverages the audio analysis capabilities of GPT-4o. The second strategy introduces a more advanced approach that uses Whisper as an audio-to-text module and evaluates the naturalness of the generated text through targeted prompt engineering, so the model is called GPT-Whisper. To the best of our knowledge, this is the first attempt to leverage ChatGPT for speech assessment. Experimental results confirm that GPT-4o alone is insufficient for accurate speech assessment. In contrast, GPT-Whisper demonstrates potential as a zero-shot speech assessment method, with yielding moderate correlation with human-based quality assessment (SRCC of 0.4360), human-based intelligibility assessment (SRCC of 0.5485), and Google ASR’s CER (SRCC of 0.5415) and notably high correlation with Whisper’s CER (SRCC of 0.7784). Furthermore, in predicting Whisper’s CER, GPT-Whisper outperforms the supervised models MOS-SSL, MTI-Net (wav2vec), and MTI-Net (Whisper) in terms of SRCC (0.7784 vs. 0.7482, 0.7418, and 0.7655). This confirms the ability of GPT-Whisper as a reliable zero-shot speech assessment method and the potential of using ChatGPT for speech assessment through the GPT-Whisper method. In the future, we plan to explore more advanced prompt engineering to further leverage ChatGPT to estimate other speech assessment metrics. We also intend to investigate direct integration of GPT-Whisper with speech-processing applications to enhance their performance.